a ski lift carried Paul up a slope at rate of 6 km/h, and he skied back down parallel to the lift at 34 km/h. The round trip took 30 minutes. How far did he ski and for how long???
Answer by Alan3354(30959)

a ski lift carried Paul up a slope at rate of 6 km/h, and he skied back down parallel to the lift at 34 km/h. The round trip took 30 minutes. How far did he ski and for how long???
-----------
The average speed of the round trip is
2*6*34/(6+34) = 10.2 km/hr
------------
RT distance = 10.2*0.5 = 5.1 km
------
2.55 km skiing
t = d/r = 2.55/34
= 0.075 hr or 4.5 minutes skiing